### v1

Abstract
To determine the causal effect of a treatment using observational data, it is important to balance the covariate distributions between treated and control groups. However, achieving balance can be difficult when treated and control groups lack overlap. In the presence of limited overlap, researchers typically choose between two types of methods: 1) methods (e.g., inverse propensity score weighting) that imply traditional estimands (e.g., ATE) but whose estimators are at risk of variance inflation and considerable statistical bias; and 2) methods (e.g., overlap weighting) which imply a different estimand, thereby changing the target population to reduce variance. In this work, we introduce a framework for characterizing estimands by their target populations and the statistical performance of their estimators. We introduce a bias decomposition that encapsulates bias due to 1) the statistical bias of the estimator; and 2) estimand mismatch, i.e., deviation from the population of interest. We propose a design-based estimand selection procedure that helps navigate the tradeoff between these two sources of bias and variance of the resulting estimators. Our procedure allows the analyst to incorporate their domain-specific preference for preservation of the original population versus reduction of statistical bias. We demonstrate how to select an estimand based on these preferences by applying our framework to a right heart catheterization study.
Keywords: average treatment effect, causal inference, inverse probability weighting, propensity score, target population
* barna126@umn.edu
Introduction
A primary goal of healthcare research is to determine the causal effect of a policy or treatment on health outcomes. Assessing causal effects requires an unconfounded comparison of treated and control groups such that the two groups have similar, or balanced, characteristics. In cases where it is not possible or ethical to balance groups through randomization, observational data must be used to assess the effectiveness of treatment. In observational data, an individuals’ covariates often influence the treatment decision which results in imbalance of the treated and control covariate distributions and, without adjustment, treatment effect estimates with the potential for considerable bias. Recent research has produced a wide variety of methods to adjust for such an imbalance. However, there can be a lack of overlap in some parts of the distribution of covariates, making it impossible to fully mitigate imbalance, resulting in large estimator bias and variance. Alternate methods aim to reduce estimator variance, through either achieving only approximate balance or by changing the population of interest such that the new population has sufficient overlap to balance the treated and control groups. Yet, these approaches, respectively, increase estimator bias or shift the target estimand away from the originally-targeted estimand. In this paper, we develop a framework for navigating the tradeoffs between preserving the original estimand while minimizing the bias and variance that can result from imbalance.
Weighting methods for balancing covariates
There are many methods for balancing covariates between groups (e.g., regression, matching, weighting), however, we will focus on weighting methods. Weighting methods reweight the data in order to achieve covariate balance between the treated and control groups. One set of such methods seek to directly balance the covariates by deriving weights subject to covariate constraints using optimization procedures (Hainmueller, 2012; Zubizarreta, 2015; Wang and Zubizarreta, 2020; Huling and Mak, 2024). Another set of methods rely on inverse-probability weighting (IPW), where each individual in the data is inverse weighted by the sample probability of receiving their assigned treatment (Rosenbaum and Rubin, 1983; Hahn, 1998; Robins, Hernán, and Brumback, 2000; Hirano and Imbens, 2001; Hirano, Imbens, and Ridder, 2003; Imbens, 2004). These methods are model-based as the propensity score, the probability of being assigned treatment, must be estimated from observed covariates to derive the weights. Both types of weights can be utilized to target different estimands defined over a specific population; for example, standard IPW weights correspond to the average treatment effect (ATE) which is defined over the total sample population. Other estimands that are commonly targeted are the average treatment effect on the treated (ATT) and on the control (ATC).
Weighting methods are popular and imply estimands with meaningful populations. Yet, when the treated and control covariate distributions lack overlap, achieving balance through these methods can be unattainable, resulting in estimator bias. When there is a lack of overlap, the use of either type of weighting methods can result in very large weights for some observations; large weights yield estimators with inflated variance and the potential for considerable bias (Busso, DiNardo, and McCrary, 2014; Hong, Leung, and Li, 2020). Direct balancing methods often provide more stable and smaller weights than propensity score methods; however, using these methods when there is a lack of overlap can still result in estimators with large variance, and in some cases, no solutions to the optimization problem (Hainmueller, 2012; Chattopadhyay, Hase, and Zubizarreta, 2020). Similar issues arise with matching methods, where large differences between the treated and control groups can result in potential bias in the treatment effect estimator due to imbalance. Thus, employing these methods for covariate balancing when there is lack of overlap produces an estimand with an interpretable and/or meaningful population of interest but whose estimator may have substantial bias and variance.
Weighting methods that address lack of overlap
A variety of new weighting methods have been created to mitigate the impacts of lack of overlap on causal estimators. For the majority of these methods, reducing the variance of the estimator is the primary goal. Within this set of methods there are two types: 1) estimator variance is reduced through increasing estimator bias; and 2) estimator variance is reduced through changing the estimand of interest such that the corresponding population has better overlap. Within this first type of method, Wang and Zubizarreta (2020) explore a class of direct balancing weights called minimal weights that aim to minimize variance when there is a lack of overlap. Minimal weights provide approximate balance, rather than exact balance of the covariate distributions, which tends to increases the bias of the estimator while reducing the variance. While estimator bias may be increased with this method, the estimand and therefore the target population remains unchanged.
Most new methods fall into the second category of methods that modify the estimand, and consequently the target population, either explicitly or implicitly. Crump et al. (2009) formalize modifying the estimand through propensity score trimming while Li and Greene (2013) and Yang and Ding (2018) propose continuous weights that approximate pair matching and propensity score trimming, respectively One of the more popular methods is overlap weights (Li, Morgan, and Zaslavsky, 2018; Li, Thomas, and Li, 2019) (with a corresponding estimand of the ATO), which weight each individual by the probability of being assigned to the opposite group. Overlap weights emphasize individuals with propensity scores around $0.5$, i.e. those with so-called clinical equipoise. When the variance of outcomes is constant, overlap weights minimize the asymptotic variance of the weighted average treatment effect estimator across all possible estimands characterized by weighted populations. Simulation studies have shown that overlap weights perform well under model misspecification and can improve statistical power compared to IPW weights (Zhou, Matsouaka, and Thomas, 2020; Mao, Li, and Greene, 2019). Of note, matching methods that address lack of overlap also change the population of interest by identifying an optimal matched population; however, the primary goal of these methods is to minimize estimator bias, rather than variance, through achieving balance (Rosenbaum, 2012; de los Angeles Resa and Zubizarreta, 2016; Visconti and Zubizarreta, 2018). For all these methods, the estimand of interest, and therefore the population, is changed in order to achieve better statistical performance of the estimator.
There are a variety of scenarios in which it is reasonable to deviate from estimands that aggregate over the entire sample population; for example, some sample populations are a convenience sample rather than a meaningful population (Rosenbaum, 2012) and as such, the ATE over such a population may not have clinical relevance. However, in many other scenarios the sample population or treated population is meaningful, making it still desirable to estimate the ATE or ATT even in the presence of low overlap. Li, Morgan, and Zaslavsky (2018) and Mao, Li, and Greene (2019) both argue that overlap weights target a population of interest; people who have propensity scores around $0.5$ may be the individuals for whom clinical treatment guidelines are unclear. While this is often true, in practice treatments may not be assigned equitably along demographic factors (Faigle et al. (2017); Raeisi-Giglou et al. (2022); Tchikrizov et al. (2023)); in these cases, it is unclear whether individuals at equipoise in the data truly reflect a group of people for whom clinical guidelines are unknown. Regardless of whether clinical equipoise is of interest, it is also often hard to describe the exact population that is induced by overlap weights. Within methods that address lack of overlap, there has been limited development on constructing an estimand that also has an interpretable population. Cardinality matching provides an option for constraining the matched set to be similar to some population of interest (Visconti and Zubizarreta, 2018). Traskin and Small (2011) and Fogarty et al. (2016) both propose methods for selecting an interpretable population with better overlap, through tree and discrete optimization techniques respectively, though both methods can still produce estimators with high variance.
Proposed framework
There are clear tradeoffs between retaining an estimand with a meaningful target population, estimator bias, and variance among current methods for treatment effect estimation when there is a lack of overlap. Standard weighting methods target estimands with meaningful populations, however there is risk of estimator bias and inflated variance. Methods such as minimal weights and overlap weights aim to either derive a set of weights or identify an estimand that minimizes estimator variance, but in doing so they sacrifice estimator bias or eschew the original target population. Of note, most methods described above aim to improve or maintain only one of estimator bias, variance, or the target population; there are few, if any, tools for the quantifying the extent with which the other two characteristics are influenced. Different applications may prioritize these three characteristics differently; in some cases statistical performance may be the primary priority, while in others it may be important to maintain the population of interest. However, as far as we know, there are few or no methods that provide tools to flexibly explore and quantify the tradeoffs between all three characteristics.
In this work, we present a conceptual framework and design-based procedure for characterizing and selecting an estimand according to its target population, estimator bias, and variance, allowing the analyst to explicitly prioritize any of these properties. As a part of the conceptual framework, we introduce a bias decomposition that encapsulates bias due to 1) the statistical bias of the estimator; and 2) estimand mismatch i.e., deviation from the population of interest. We present a set of estimands from which to select an optimal estimand, where the ATE, ATT, ATC, and ATO are special cases of the set. To characterize this set of estimands, we propose two design-based metrics that target each part of the bias decomposition. Furthermore, we propose an estimand selection procedure that clearly illustrates the bias-variance tradeoffs of potential estimands. We analyze synthetic data from a variety of data generating scenarios and data from a right heart catheterization study to demonstrate the effectiveness of our method in navigating these tradeoffs. We present all methods and results in reference to the ATE population as the population of interest, however, all methods can apply to other estimands/populations such as the ATT or ATC.
The remainder of the paper is outlined as follows. In Section 2 we introduce notation and our conceptual framework for characterizing and selecting estimands. Section 3 describes the design-based metrics for characterizing both types bias and the proposed estimand selection procedure. Section 4 validates our metrics and evaluates the performance of our estimation selection procedure in comparison to the ATE and ATO through simulation studies. In Section 5 we apply our entire procedure to reanalyze a study on right heart catheterization. Finally, Section 6 discusses the results and future work.
Conceptual framework for causal estimand selection
Notation and assumptions
Consider a sample $\{(Z_{i},Y_{i},\bm{X}_{i})\}_{i=1}^{n}$ of size $n$ from a population. For individual or unit $i$, let $Z_{i}=z$, indicate belonging to the treatment ($z=1$) or control group ($z=0$), $Y_{i}$ be the outcome, and $\bm{X}_{i}=(X_{i1},\ldots X_{ip})$ be a $p$ length vector of covariates. Let $e(\bm{x})=\Pr(Z_{i}=1|\bm{X}_{i}=\bm{x})$ be the propensity score. We use the potential outcomes framework (Neyman (1990); Rubin (1974, 1978); Hernan and Robins (2024)), where $Y(z)$ is the outcome that would be observed if assigned to treatment group $z$. In practice, only one potential outcome is observed per individual. We assume the standard stable unit treatment value assumption (SUTVA) that states the potential outcomes for each individual are unaffected by the treatment assignments of other individuals and that there is only one version of treatment, which implies that $Y_{i}=Y_{i}(Z_{i})$. We also assume that the treatment effect is strongly unconfounded such that $\{Y(0),Y(1)\}{\,\perp\!\!\!\perp\,}Z|\bm{X}$, which requires that there are no unmeasured confounders. Finally, we assume positivity (Rosenbaum and Rubin (1983)), which requres that $10<e(\bm{x})<$ 1 such that all individuals have a non-zero probability of being assigned to either the treatment or control group. Given these assumptions and using potential outcomes notation, the conditional average treatment effect (CATE) is $\tau(\bm{x})=\mu_{1}(\bm{x})-\mu_{0}(\bm{x})=E\{Y(1)-Y(0)|\bm{X}=\bm{x}\}$. Often, we desire to compute the ATE not with respect to a single $\bm{x}$ but over a specific target population or covariate distribution. Let $f(\bm{x})$ be the marginal distribution of the covariates and $h(\bm{x})$ some function such that $f(\bm{x})h(\bm{x})$ is the density of a target population. Then the estimand, $\tau_{h}$, averages the CATE over the target population,
[…]
Note that each unique $h(\bm{x})$ defines a different target population and thus a different causal estimand, $\tau_{h}$. Throughout, we assume that $\tau(\bm{x})$ can vary with $\bm{x}$, such that averaging over different populations (i.e., different causal estimands) may yield different results. For a given weight function $h(\bm{x})$, Li, Morgan, and Zaslavsky (2018) propose a class of propensity score balancing weights such that weighted covariate distributions are balanced between the two treatment groups,
$w_{1}(\bm{x})$ $=\frac{h(\bm{x})}{e(\bm{x})}\>\>\>\text{for}\>\>\>Z=1,\text{ and}$ (1)
$w_{0}(\bm{x})$ $=\frac{h(\bm{x})}{1-e(\bm{x})}\>\>\>\text{for}\>\>\>Z=0.$ (2)
One common causal estimand is the ATE, which averages $\tau(\bm{x})$ over the entire covariate population from which the sample was derived. For the ATE, $h(\bm{x})=1$ and the balancing weights are the IPW weights. Other common average casual estimands and their corresponding $h(\bm{x})$, weights, and target populations are in Table 1. For a given $h(\bm{x})$, the sample estimator for $\tau_{h}$ is
$\hat{\tau}_{h}=\frac{\sum_{i}w_{1}(\bm{X}_{i})Z_{i}Y_{i}}{\sum_{i}w_{1}(\bm{X}_{i})Z_{i}}-\frac{\sum_{i}w_{0}(\bm{X}_{i})(1-Z_{i})Y_{i}}{\sum_{i}w_{0}(\bm{X}_{i})(1-Z_{i})},$
where $\hat{\tau}_{h}$ is consistent for $\tau_{h}$ when the weights in Equations (1) and (2) are used (Li, Morgan, and Zaslavsky, 2018). For the following sections, aside from Section 2.4, we generally assume $\hat{\tau}_{h}$ is computed with estimated propensity scores as the the true propensity scores are rarely known.
Table 1: Causal estimands and their corresponding $h(\bm{x})$, balancing weights (Li, Morgan, and Zaslavsky, 2018), and population.
Estimand $h(\bm{x})$ $w_{0}(\bm{x}),w_{1}(\bm{x})$ Target Population
ATE $1$ $1/\{1-e(\bm{x})\},1/e(\bm{x})$ Combined
ATT $e(\bm{x})$ $e(\bm{x})/\{1-e(\bm{x})\},1$ Treated
ATC $1-e(\bm{x})$ $1,e(\bm{x})/\{1-e(\bm{x})\}$ Control
ATO $e(\bm{x})\{1-e(\bm{x})\}$ $e(\bm{x}),1-e(\bm{x})$ Overlap
Decomposing sources of error in the presence of positivity violations
In practice, one may need to mitigate variability and/or bias due to lack of overlap by modifying the estimand from the original estimand. We call the modified estimand a “performance” estimand, as it is often chosen pragmatically. In this section we aim to characterize the sources of discrepancy between the original target estimand of interest and the specific estimator used in practice, which may or may not target a performance estimand. This characterization allows us to systematically navigate the bias and variance tradeoffs that account for both statistical errors and mismatch between the targeted estimand and performance estimand.
Population target estimands are estimands such that $h(\bm{x})$ induces a meaningful population of interest and would typically be viewed as the key estimand of scientific interest. Performance estimands, on the other hand, are often chosen to provide improved statistical benefits (e.g. lower estimator bias, variance) in comparison to the population target estimand. The population target and performance estimands can be the same; for example, the ATO provides variance reduction compared to other estimands and may also correspond to a population of interest. However, we will generally consider estimands such as the ATO and those induced by propensity score trimming as performance estimands as these estimands were originally proposed in order to achieve estimators with reduced variance. Throughout the rest of the paper, we will consider the ATE ($h(\bm{x})=1$), $\tau$, as the population target estimand for demonstrative purposes, though the proposed methods and framing can straightforwardly apply to any target estimand.
Let $h_{p}(\bm{x})$ correspond to a performance estimand $\tau_{h_{p}}$ for a given analysis where the target estimand is $\tau$. Then the bias of $\hat{\tau}_{h_{p}}$ in relation to $\tau$ can be decomposed as
$E(\hat{\tau}_{h_{p}})-\tau\>\>=\>\>\underbrace{E(\hat{\tau}_{h_{p}})-\tau_{h_{p}}}_{\text{Statistical bias}}\>\>\>+\!\!\!\!\underbrace{\tau_{h_{p}}-\tau}_{\text{Estimand mismatch}}.$ (3)
The first term represents the statistical bias of the estimator $\hat{\tau}_{h_{p}}$, while the second term represents the mismatch between the ATE and the performance estimand. To minimize the discrepancy between $\hat{\tau}_{h_{p}}$ and the target estimand, both the estimand mismatch and the statistical bias of $\hat{\tau}_{h_{p}}$ must be controlled. Note that estimand mismatch and statistical bias are often at odds, especially in scenarios with substantial positivity violation; for example, $\tau_{h_{p}}=\tau$ would yield zero estimand mismatch, but $\hat{\tau}$ tends to have high statistical bias when there is low overlap. To reduce overall bias, it is desirable to identify a $\tau_{h_{p}}$ such that $\hat{\tau}_{h_{p}}$ has smaller statistical bias than $\hat{\tau}$ as well as limited estimand mismatch.
To help navigate the tradeoffs between estimand mismatch, statistical bias, and variance we propose an approach for the selection of the performance estimand, $\tau_{h_{p}}$. We allow for selection of $\tau_{h_{p}}$ from the following family of potential weight functions,
[…]
where the ATE, ATC, ATT, ATO are all special cases of this set (Figure 1). By varying the two parameters, $c$ and $d$, this family provides a smooth transition between each of these widely-used estimands. We can then select an optimal $\tau_{h_{p}}$ from this set by characterizing the bias and variance of their corresponding estimators.
Figure 1: We show the normalized $h(\bm{x})$ in relation to the propensity score, $e(\bm{x})$, for a set of estimands. For each subplot, we plot $h(\bm{x})$ for a well-known estimand and the ATO and demonstrate how estimands within our proposed set (Equation (4)) are intermediaries between them.
Identifying measurable attributes of estimand mismatch and statistical bias
To leverage the bias decomposition proposed in the previous section for estimand selection, it is necessary to understand what measurable attributes impact estimand mismatch and statistical bias. We will focus on the ATE as the target estimand, however, this work can also be applied to the other estimands of interest such as the ATT, ATC, or even when transporting the effect to a new population. For clarity in notation, we assume that the weights have been normalized such that $\sum_{i=1}^{n}w_{z}(\bm{X}_{i})I(Z_{i}=z)=n_{z}$, the sample size of treatment group $z$. We can then express the difference between a performance estimator and the ATE as
$\hat{\tau}_{h_{p}}-\tau={}$ $\underbrace{\int\mu_{1}(\bm{x})d\{F_{n_{0},0,\bm{w}_{0}}-F_{n}\}(\bm{x})-\int\mu_{0}(\bm{x})d\{F_{n_{1},1,\bm{w}_{1}}-F_{n}\}(\bm{x})}_{\text{Estimand mismatch}}$ (5)
+$+\underbrace{\int\{\mu_{1}(\bm{x})+\mu_{0}(\bm{x})\}d\{F_{n_{1},1,\bm{w}_{1}}-F_{n_{0},0,\bm{w}_{0}}\}(\bm{x})}_{\text{Statistical bias}}$ (6)
$-\int\{\mu_{1}(\bm{x})-\mu_{0}(\bm{x})\}d\{F-F_{n}\}(\bm{x})$ (7)
$+\frac{1}{n_{1}}\sum_{i=1}^{n}w_{1}(\bm{X}_{i})\epsilon_{i}Z_{i}-\frac{1}{n_{0}}\sum_{i=1}^{n}w_{0}(\bm{X}_{i})\epsilon_{i}(1-Z_{i}),$ (8)
where $\epsilon_{i}=Y_{i}(Z_{i})-\mu_{Z_{i}}(\bm{X}_{i})$, $F_{n}(\bm{x})=\sum_{i=1}^{n}I(\bm{X}_{i}\leq\bm{x})/n$ is the empirical CDF, and $F_{n_{z},z,\bm{w}_{z}}(\bm{x})=\sum_{i=1}^{n}w_{z}(\bm{X}_{i})I(\bm{X}_{i}\leq\bm{x},Z_{i}=z)/n_{z}$ is the weighted empirical CDF. In this expression, term (7) goes to zero for a representative sample of the population and term (8) has mean zero. Then, the difference between our performance estimator and population target estimand can be characterized by terms (5) and (6) which are induced by estimand mismatch and statistical bias, respectively. We note that term (5) often has non-zero expectation even with no estimand mismatch (the expectation is zero when using the true propensity scores); however, this term does tend to increase in size as estimand mismatch increases. The measurable quantities in terms (5) and (6) are 1) the imbalance between the empirical CDFs $F_{n}$ and $F_{n_{z},z,\bm{w}_{z}}$, and 2) the imbalance between empirical CDFs $F_{n,0,\bm{w}_{0}}$ and $F_{n,1,\bm{w}_{1}}$. To formally characterize estimand mismatch and statistical bias for a given $\hat{\tau}_{h_{p}}$, one can then quantify the distance between empirical CDFs $F_{n}$ and $F_{n_{z},z,\bm{w}_{z}}$ and empirical CDFs $F_{n_{0},0,\bm{w}_{0}}$ and $F_{n_{1},1,\bm{w}_{1}}$. We propose design-based metrics that measure these two distances in Sections 3.1 and 3.2. To select an optimal performance estimand from the family of $h(\bm{x})$ in Equation (4), we evaluate each $h_{c,d}(\bm{x})$ by these two metrics.
Identifying estimands with minimal estimator variance
While we have focused on estimator bias, we also often want to identify a performance estimand whose corresponding estimator has small or minimal variance. Li, Morgan, and Zaslavsky (2018) prove that overlap weights ($h(\bm{x})=e(\bm{x})\{1-e(\bm{x})\}$) yield the estimator with the minimum asymptotic variance for all $\hat{\tau}_{h}$ under homoscedasticity. However, overlap weights do not necessarily yield the minimum asymptotic variance when homoscedasticity does not hold. The following proposition establishes the variance conditions under which $\hat{\tau}_{h*}$ has minimum asymptotic variance (proof in Supplementary Materials).
Proposition 1 Let $v_{z}(\bm{x})=Var\{Y(z)|\bm{X}=\bm{x}\}$ and $h^{*}(\bm{x})>0$. Then if
$v_{0}(\bm{x})$ $=v\frac{k_{0}(\bm{x})\{1-e(\bm{x})\}}{h^{*}(\bm{x})\{k_{0}(\bm{x})+k_{1}(\bm{x})\}}\text{ and}$
$v_{1}(\bm{x})$ $=v\frac{k_{1}(\bm{x})e(\bm{x})}{h^{*}(\bm{x})\{k_{0}(\bm{x})+k_{1}(\bm{x})\}},$
$\hat{\tau}_{h^{*}}$ has the minimum asymptotic variance of all balancing weight estimators $\hat{\tau}_{h}$ for $v\in\mathbb{R}^{+}$ and functions $k_{0}(\bm{x}),k_{1}(\bm{x})>0$ that are well defined over the domain of $\bm{X}$.
This result further motivates selecting a performance estimand specific to a given dataset, as the overlap weights estimator may not have the minimum asymptotic variance let alone the minimum finite sample variance for some data. In order to characterize estimator variance and select an optimal performance estimand from the family of $h(\bm{x})$ in Equation (4), we estimate the standard error of each $\hat{\tau}_{h_{c,d}}$ using a bootstrapped standard error estimator (Crump et al., 2009).
Implementation of causal estimand selection framework
Characterizing estimand mismatch
We use the weighted energy distance as our measure of imbalance between distributions (Huling and Mak, 2024). We choose the energy distance due to its performance in high dimensions and computational efficiency, however other measures of distributional imbalance such as the maximum mean discrepancy, Wasserstein distance, etc. could work well and be used here instead. The weighted energy distance between empirical CDFs $F_{n}$ and $F_{n_{z},z,\bm{w}_{z}}$ is
$\mathcal{E}(F_{n_{z},z,\bm{w}_{z}},F_{n})=\frac{2}{n_{z}n}\sum_{i=1}^{n}\sum_{j=1}^{n}w_{z}(\bm{X}_{i})I(Z_{i}=z)||\bm{X}_{i}-\bm{X}_{j}||_{2}$
$-\frac{1}{n_{z}^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}w_{z}(\bm{X}_{i})w_{z}(\bm{X}_{j})I(Z_{i}=Z_{j}=z)||\bm{X}_{i}-\bm{X}_{j}||_{2}-\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}||\bm{X}_{i}-\bm{X}_{j}||_{2}.$
In this context, $\mathcal{E}(F_{n_{z},z,\bm{w}_{z}},F_{n})$ quantifies the difference between the population that is generated by $h_{c,d}(\bm{x})$ and the target ATE population. While non-zero energy distances indicate the degree of difference between two distributions, the energy distance is not a standardized metric and therefore is not comparable across different weight functions. To create a metric comparable across all $h_{c,d}(\bm{x})$, we construct a permutation-based statistical test for $H_{0}:F=F_{z,\bm{w}_{z}}$. The proposed test procedure for a given $F_{n}$ and $F_{n_{z},z,w_{z}}$ follows:
  1. 1.
    Compute the test statistic $\mathcal{E}(F_{n_{z},z,\bm{w}_{z}},F_{n})$
  2. 2.
    For $b=1,\ldots,B$
    • Sample $F^{(b)}_{n_{z}}$, of size $n_{z}$ from $F_{n}$ without replacement
    • Compute $\mathcal{E}(F^{(b)}_{n_{z},\bm{w}_{z}},F_{n})$
  3. 3.
    Compute the $p$-value, $p=\frac{1}{B}\sum_{b=1}^{B}I[\mathcal{E}(F_{n_{z},z,\bm{w}_{z}},F_{n})\leq\mathcal{E}(F^{(b)}_{n_{z},\bm{w}_{z}},F_{n})]$
In step 2, since $\bm{w}_{z}$ has no relationship to the randomly sampled $\{F^{(b)}_{n_{z}}\}_{b=1}^{B}$, $\{F^{(b)}_{n_{z},\bm{w}_{z}}\}_{b=1}^{B}$ approximates $F_{z,\bm{w}_{z}}$ under the desired null $F=F_{z,\bm{w}_{z}}$. We can use this $p$-value for inference or simply as a standardized measure of the imbalance between $F_{n_{z},z,\bm{w}_{z}}$ and $F_{n}$, where smaller $p$-values tend to correspond to $\hat{\tau}_{h_{c,d}}$ with the potential for larger estimand mismatch. If we want to determine whether the population induced by $h_{c,d}(\bm{x})$ is different than the ATE population, we reject the null hypothesis that these populations are the same at $p<\alpha=0.05$.
Characterizing statistical bias
Let $G$ indicate the CDF for the fitted propensity scores $\hat{e}(\bm{X_{i}})$. The weighted energy distance between empirical CDFs $G_{n_{0},0,\bm{w}_{0}}$ and $G_{n_{1},1,\bm{w}_{1}}$ is
$\mathcal{E}(G_{n_{0},0,\bm{w}_{0}},G_{n_{1},1,\bm{w}_{1}})=\frac{2}{n_{0}n_{1}}\sum_{i=1}^{n}\sum_{j=1}^{n}w_{0}\{\hat{e}(\bm{X}_{i})\}w_{1}\{\hat{e}(\bm{X}_{j})\}I(Z_{i}=0)I(Z_{j}=1)||\hat{e}(\bm{X}_{i})-\hat{e}(\bm{X}_{j})||_{2}$
$-\sum_{z=0}^{1}\frac{1}{n_{z}^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}w_{z}\{\hat{e}(\bm{X}_{i})\}w_{z}\{\hat{e}(\bm{X}_{j})\}I(Z_{i}=Z_{j}=z)||\hat{e}(\bm{X}_{i})-\hat{e}(\bm{X}_{j})||_{2}.\vspace{-0.1in}$
We use the propensity score empirical CDFs instead of the covariate empirical CDFs (as seen in term (6)) because the use of propensities scores has been shown to be sufficient for covariate sample balance (Rosenbaum and Rubin, 1983) and to simplify computation. The permutation-based statistical test procedure for $H_{0}:G_{0,\bm{w}_{0}}=G_{1,\bm{w}_{1}}$ follows:
  1. 1.
    Compute $\mathcal{E}(G_{n_{0},0,\bm{w}_{0}},G_{n_{1},1,\bm{w}_{1}})$
  2. 2.
    For $b=1,\ldots,B$
    • Permute the treatment assignments to generate $G^{(b)}_{n_{0}}$ and $G^{(b)}_{n_{1}}$
    • Compute $\mathcal{E}(G^{(b)}_{n_{0}},G^{(b)}_{n_{1}})$
  3. 3.
    $p=\frac{1}{B}\sum_{b=1}^{B}I[\mathcal{E}(G_{n_{0},0,\bm{w}_{0}},G_{n_{1},1,\bm{w}_{1}})\leq\mathcal{E}(G^{(b)}_{n_{0}},G^{(b)}_{n_{1}})]$
In step 2, permuting the treatment assignment breaks the relationship between treatment assignment and covariates (and therefore $\bm{w}_{z}$) such that $\{G_{n_{0}}^{(b)},G_{n_{1}}^{(b)}\}_{b=1}^{B}$ approximate $G_{0},G_{1}$ under the desired null $G_{0,\bm{w}_{0}}=G_{1,\bm{w}_{1}}$. Here, we primarily use $p$ as a standardized measure of the imbalance between $G_{n_{0},0,\bm{w}_{0}}$ and $G_{n_{1},1,\bm{w}_{1}}$, where smaller $p$-values tend to correspond to $\hat{\tau}_{h_{c,d}}$ with the potential for larger statistical bias. We are less interested in specifying an $\alpha$ level as $\bm{w}_{0}$, $\bm{w}_{1}$ are designed to make these distributions equal and thus we expect these $p$-values to generally be high. While this $p$-value is motivated by characterizing statistical bias due to lack of overlap, in practice it characterizes statistical bias due to any source(s), such as propensity model misspecification.
Estimand selection
There are many potential ways to select an optimal estimand for a dataset given the three metrics described in the previous sections. We propose the following selection procedure as an approach for enumerating a sequence of estimators, each approximately optimal for a given subjective choice in emphasis of estimand mismatch versus statistical bias. Thus, the procedure allows the analyst to choose a point along this spectrum most appropriate for their application. The procedure is as follows:
  1. 1.
    Compute $p$ as defined in Sections 3.1 and 3.2 and the bootstrapped standard error estimator for some grid of potential $h_{c,d}(\bm{x})$, $c,d\in[0,1]$ (in Section 4 we use all combinations of $c,d=0,0.05,0.10,\ldots,1$).
  2. 2.
    For each set of $p$-values (i.e., corresponding to estimand mismatch and statistical bias) create density contours on the $(c,d)\in[0,1]\times[0,1]$ space (Figure 2a). Interpolation may be needed to smooth the contours. Next, use kriging to interpolate the estimated standard errors across a finer grid over $[0,1]\times[0,1]$.
    Figure 2: Subfigure a) shows the contours of the $p$-values corresponding to estimand mismatch and statistical bias. Data is interpolated using cubic splines in order to smooth contours. Subfigure b) shows the intersection of each estimand mismatch $p$-value contour with the largest statistical bias $p$-value contour level (color indicates the estimand mismatch $p$-value contour). The points in subfigure b) show the optimal estimands for each estimand mismatch $p$-value contour.
  3. 3.
    Intersect the estimand mismatch and statistical bias $p$-value contours. For each estimand mismatch $p$-value contour level, indentify the area intersecting with the largest statistical bias $p$-value contour level and select the $h_{c,d}$ with the smallest estimated standard error in that area (Figure 2b).
The first step in the procedure characterizes a grid of estimands by our estimand mismatch and statistical bias metrics as well as estimated standard error; finer grids take longer to compute but provide more complete information. The second step in this process simply extends this characterization across either all estimands defined by Equation (4) or a finer grid of estimands. Step three identifies and selects an estimand with low potential for statistical bias and the smallest estimated standard error within each estimand mismatch $p$-value contour. This results in a sequence of optimal (i.e., low potential for statistical bias and low variance) estimands; as the estimand mismatch $p$-value contour level gets smaller, the selected estimand tends to have larger estimand mismatch but smaller estimator statistical bias and variance.
In general, we recommend selecting the optimal estimand within the the $(0.05,0.1]$ $p$-value contour, as this estimand will tend to have the lowest standard error among estimands where we fail to reject the null hypothesis that the weighted estimand population is the same as the ATE population. Although we propose this procedure and implement it within Sections 4 and 5, it can be varied depending on the particular data application. For example, if only the statistical performance of the estimator is of interest, one could select the estimand in the largest statistical bias $p$-value contour with the smallest estimated standard error. Ultimately, the selection procedure can be adapted based on the perceived importance of estimand mismatch, statistical bias, and estimator variance in a given application.
Simulation
Methods
We construct data generating scenarios with varying levels of propensity score overlap, treatment effect heterogeneity, and percent of treated individuals. Through these scenarios, we first aim to validate the metrics proposed in Sections 3.1 and 3.2 across a variety of data generating processes. We further aim to determine whether the ability of these metrics to characterize estimand mismatch and statistical bias varies across data generating scenarios. In addition, we aim to illustrate whether and when our estimand selection procedure can identify estimands whose corresponding estimators have benefits over the ATE and ATO estimators. By exploring scenarios with different propensity score overlap and treatment effect heterogeneity, we also demonstrate how these attributes influence the bias-variance tradeoffs of estimand selection within our procedure.
Data generation and characterizing estimands
We implement nearly the same simulation data generation process as in Li, Thomas, and Li (2019). Here, we drop the observation index on variables for clarity of presentation. We generate $V_{1},\ldots,V_{6}$ from a multivariate normal distribution with $E[V_{i}]=0$, $Var(V_{i})=1$ for all $i=1,\ldots,6$ and $Cov(V_{i},V_{j})=0.5$ for all $i\neq j$. We take $X_{1}\textrm{--}X_{3}=V_{1}\textrm{--}V_{3}$ and dichotomize $V_{4}\textrm{--}V_{6}$ such that $X_{j}=I[V_{j}<0]$ for $j=4,5,6$. The true propensity scores are generated as $e(\bm{X})=\{1+\exp(-\alpha_{0}-\sum_{j=1}^{6}\alpha_{j}X_{j})\}^{-1}.$ Our continuous outcome, $Y$, is generated as $Y\sim N\{E(Y|Z,\bm{X}),1\}$ where $E(Y|Z,\bm{X})=\beta_{0}+\sum_{j=1}^{6}\beta_{j}X_{j}+\Delta(\bm{X})Z$. Similar to Mao, Li, and Greene (2019), we choose two heterogeneous treatment effect functions, $\Delta(\bm{x})$, that have a concave relationship with the true propensity scores:
$\Delta(\bm{x})$ $=8e(\bm{x})\{1-e(\bm{x})\}$ (9)
$\Delta(\bm{x})$ $=16e(\bm{x})\{1-e(\bm{x})\}-1.$ (10)
Both functions are centered at $1$, but Equation (9) has range $(0,2)$ while Equation (10) has range $(-1,3)$; therefore, these functions represent a medium and high treatment effect heterogeneity, respectively.
Consistent with Li, Thomas, and Li (2019), the parameters for the true propensity score model are $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{6})=(0.15\gamma,0.3\gamma,0.3\gamma,-0.2\gamma,-0.25\gamma,-0.25\gamma)$. Throughout the simulations we explore $\gamma=1,2,3,4$ where $\gamma=1$ generates data with substantial propensity score overlap and $\gamma=4$ generates data with minimal propensity score overlap and substantial density of propensity scores in the tails (Supplementary Figure 1). We choose $\alpha_{0}$ to generate the desired proportion of treated individuals in the data ($0.25$ and $0.5$). For our outcome model, we set $(\beta_{0},\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5},\beta_{6})=(0,-0.5,-0.5,-1.5,0.8,0.8,1)$.
We have a total of 16 simulaion scenarios that represent all combinations of $\gamma=1,2,3,4$, $25\%$ and $50\%$ treated individuals, and medium and high treatment effect heterogeneity. For each simulation scenario we generate 1000 independent datasets with 1000 observations each. For each simulated dataset we explore a grid of candidate estimand values over $(c,d)\in[0,1]\times[0,1]$. We use all combinations of $c,d=0,0.05,0.1,\ldots,1$ for a total of $441$ estimands. We calculate the true value of each estimand using 10 million Monte Carlo samples. For each estimand, we calculate the average treatment effect estimate, the absolute percent statistical bias and average estimated standard error of the estimator, empirical standard deviation, and the average $p$-values from the two proposed measures of distributional imbalance (Sections 3.1 and 3.2). For our measure of distributional imbalance between $F_{n}$ and $F_{n,z,w}$, we calculate two $p$-values, one comparing each weighted treatment group to the targeted population. For all simulation results, we use the minimum of these two $p$-values as the main design-based metric for characterizing estimand mismatch.
Estimand selection
Using the results from 4.1.1, we select a sequence of optimal estimands per 1000 datasets in each data generating scenario using our proposed estimand selection procedure (Section 3.3). We use cubic spline interpolation to smooth the $p$-value contours and interpolate standard error estimates using kriging over a $500\times 500$ grid with a Gaussian variogram model. For each dataset, we select an optimal estimand for each estimand mismatch $p$-value contour. However, this results in a different number of optimal estimands per dataset; for example, one dataset may have estimand mismatch $p$-values ranging from 0 - 0.7 (set of 16 optimal estimands) and another may have $p$-values only ranging from 0 - 0.3 (set of 8 optimal estimands). Therefore, for each estimand mismatch $p$-value contour (besides $[0,0.025]$) only a subset of the 1000 datasets have a corresponding optimal estimand; the larger the $p$-value contour, the fewer datasets that have an optimal estimand within that contour. Since the set of datasets that have a specific range of estimand mismatch $p$-values may differ systematically, we select two different sample size cutoffs ($n=500,900$) to compute aggregate metrics over. Effectively, we 1) determine the largest $p$-value contour, $p^{*}$, with at least $n$ datasets with optimal estimands; 2) identify the specific datasets with optimal estimands in the $p^{*}$ contour; and 3) compute aggregate metrics for all $p$-value contours with $p<p^{*}$ over the datasets identified in step 2). We calculate the MSE with respect to the true ATE for the ATE estimator, the estimators corresponding to estimands selected with our method, and the ATO estimator.
Results
Characterizing estimand mismatch
For the 50% treated scenarios, the estimands with $p>0.05$ (i.e., the estimands whose weighted covariate populations are similar to the ATE covariate population) follow the region along the diagonal ($c=d$) which, at high propensity score overlap, goes between the ATE to the ATO (Figure 3, left panel). For all of these estimands, $h_{c,d}(\bm{x})$ peaks around propensity scores of $0.5$. In contrast, for the 25% treated scenarios, the estimands with $p>0.05$ follow a trend above the diagonal ($c<d$) such that $h_{c,d}(\bm{x})$ peaks at propensity scores around $0.25$. Therefore, the estimands that induce a covariate population similar to the ATE population are estimands such that $h_{c,d}(\bm{x})$ peaks at the proportion treated individuals within the data. As propensity score overlap decreases, fewer estimands yield $p>0.05$ and these estimands tend to have $c,d$ close to the ATE (e.g. $c,d<0.25$).
Comparing the 25% treated columns between the left and right panels in Figure 3, there is a clear relationship between the $p$-values and absolute estimand mismatch; estimand mismatch tends to be smallest along a linear trend above the diagonal. For the 50% treated scenario the relationship is not as immediately clear. Estimands close to the ATE do tend to have low estimand mismatch, however, many estimands further from the ATE along the diagonal have comparatively large estimand mismatch. Estimand mismatch depends not only on distributional imbalance, but also on how the treatment effect varies across the covariate distribution (see term (5)). Therefore, $p<0.05$ from our proposed measure of distributional imbalance indicates estimands with the potential for large absolute estimand mismatch, but the specific estimands with large estimand mismatch vary by data generating mechanism. Supplementary Figure 2 presents absolute estimand mismatch when the treatment effect has a linear relationship with the propensity score, rather than quadratic. For this data generating process with 50% treated individuals, the estimands with the smallest estimand mismatch follow the diagonal regardless of how far these estimands get from the ATE.
For both 25% treated data generating scenarios shown in Figure 3 and Supplementary Figure 2, no estimands with absolute estimand mismatch above the 0.20 quantile (within a data generating scenario) have $p>0.05$. For 50% treated scenarios, no estimands with an absolute estimand mismatch above the 0.45 quantile have $p>0.05$, except for the $\gamma=1,2$ scenarios from Figure 3. However, in these two scenarios the absolute estimand mismatch is relatively small and therefore selecting any of the proposed estimands will yield low estimand mismatch. Thus, the $p$-value metric protects against selecting estimands with high estimand mismatch across a variety of data generating mechanisms.
Figure 3: The left panel presents the average of the minimum of the two estimand mismatch related $p$-values, calculated as described in Section 3.1, for each estimand. The right panel presents the absolute difference between the true estimand values and the true ATE (i.e., the absolute estimand mismatch) under the medium $\Delta$ heterogeneity scenario. Columns distinguish the 25% and 50% treated scenarios while rows distinguish the different propensity score overlap scenarios.
Characterizing statistical bias
For both proportion treated scenarios, the $p$-values characterizing statistical bias tend to be largest at $c,d>0.75$ (Figure 4, left panel). As propensity score overlap decreases the $p$-values get smaller, indicating that the estimators have a higher potential for statistical bias. For the 50% treated scenario, the region of higher $p$-values tends to move from the ATO to ATE along the diagonal, with a majority of these estimands being closer to ATO. While most of the higher $p$-value estimands are also close to the ATO for the 25% treated scenario, estimands above the diagonal ($c<d$) tend to have lower $p$-values (and therefore a higher potential for statistical bias) relative to the 50% treated scenario.
[…]
Figure 4: The left panel presents the average $p$-value corresponding to the statistical bias metric, calculated as described in Section 3.2, for each estimand. The right panel presents the relationship and Spearman correlation coefficient ($\rho$) between $\log_{10}(p)$ and the absolute percent statistical bias of the estimator for each estimand. Columns distinguish the 25% and 50% treated scenarios while rows distinguish the different propensity score overlap scenarios.
Standard error of estimators
We simulate the data to have homoscedastic residual variance such that our results are consistent with Li, Morgan, and Zaslavsky (2018); the estimand corresponding to the minimum variance estimator tend to be close to the ATO ($c=d=1$) regardless of data generating scenario (Supplementary Figure 4). However, as shown in Proposition 1, this relationship could change when the residual variance is not homoscedastic. Of additional note, lack of propensity score overlap also leads to bias in variance estimation. When there is low propensity score overlap, the average standard error (SE) tends to underestimate the true SE of the estimators (as measured by the empirical standard deviation (SD)) (Supplementary Figure 5). Mao, Li, and Greene (2019) found the same relationship when using the sandwich variance estimator in this setting. Despite this discrepancy, there is a very strong relationship between the relative ranking of the average estimated SE and empirical SD (all but one scenario has $\rho>0.985$) across all data generating scenarios. This indicates that the bootstrapped SE estimator is still suitable for comparing and selecting estimands with minimal or small variance for a given dataset.
Estimand selection
Across data generating scenarios (for a $n=500$ cutoff), the estimators corresponding to estimands selected with our method have a lower mean squared error (MSE) with respect to the ATE than either the ATE or ATO estimators (Figure 5); this indicates that our procedure can identify estimands that have better overall performance than simply targeting either extremal estimand. From left to right in each subplot of Figure 5, the points depict estimands that become less similar the ATE and more similar to the ATO. These curves, specifically within low-medium propensity score overlap ($\gamma=3$) scenarios, help demonstrate the bias-variance tradeoff when the scientific interest is with respect to the ATE rather than a performance estimand. For medium $\Delta$ heterogeneity, both the ATE (tends to have large variance) and ATO (tends to have substantial bias due to estimand mismatch) estimators have a large MSE indicating that bias and variance contribute relatively equally to the MSE. Thus, the estimators selected by our method all have lower MSE than the ATE and ATO estimators by moderating the size of both bias and variance. In contrast, for high $\Delta$ heterogeneity, the ATO estimator has a much larger MSE than the ATE estimator, indicating that bias (specifically estimand mismatch) dominates the MSE. Thus, estimators corresponding to larger estimand mismatch $p$-value contour levels tend to have the smallest MSE as they have reduced statistical bias and variance compared to the ATE estimator without incurring substantial estimand mismatch. When using a sample size cutoff of $n=900$, we identify similar trends among the data generating scenarios (Supplementary Figure 6). The average estimand per estimand mismatch $p$-value contour level, across all possible datasets, is shown in Supplementary Figure 7 to demonstrate on average which $c,d$ values are selected.
Figure 5: Mean squared error (MSE) with respect to the true ATE for the estimators for the ATE, the estimands selected with our procedure, and the ATO by simulation scenario. Each of the 16 subplots corresponds to a single simulation scenario as labelled by the top and right axes. Estimands selected from our framework are labeled by the lower bound of their corresponding estimand mismatch $p$-value contour. Estimands selected (i.e., estimand mismatch $p$-value contours present) in at least 500 of the simulated datasets are shown.
Consequently, there is a clear intuition for selecting a single estimand from the set of optimal estimands identified through our method. While a default estimand selection could be the estimand selected within the $(0.05,0.1]$ $p$-value contour, one can use both subject knowledge of the expected treatment effect heterogeneity and preferences with respect to variance and bias to select a single estimand. If treatment effect heterogeneity is expected to be low, the estimands from smaller $p$-value contours will tend to minimize the MSE; if treatment effect heterogeneity is expected to be high, the estimands from larger $p$-value contours will tend to minimize MSE. For applications where variance reduction has higher priority than retaining the target estimand, either the ATO or the estimands from the smallest $p-$value contours will tend to be better estimand choices. For applications where bias reduction with respect to the target estimand is the priority, estimands from larger $p$-value contours will tend to perform best. In general, estimands selected from our method tend to have lower MSE than either the ATE or ATO because they are intermediate options within the bias-variance tradeoff; they tend to mitigate the size of statistical bias, estimand mismatch, and variance rather than minimizing just one of these attributes.
Application to right heart catheterization data
Right heart catheterization (RHC) is an invasive medical procedure that is performed on critically ill patients in order to measure cardiac function. An observational study by Connors et al. (1996) examined the impact of RHC on survival for patients in the ICU. The study was conducted between 1989-1994 across five medical centers within the United States. To account for treatment selection bias within the data, the authors matched treated and untreated individuals by propensity score matching as part of their analysis (Rosenbaum and Rubin (1983, 1983)). The data has also been re-analyzed in various other works (Hirano and Imbens, 2001; Crump et al., 2009; Rosenbaum, 2012; Li, Morgan, and Zaslavsky, 2018; Huling and Mak, 2024). The primary analysis showed that RHC may be detrimental to survival, which was contrary to the popular belief that RHC improved patient outcomes.
The data consists of 5735 individuals, with 2184 treated (RHC applied within 24 hours of hospital admission) and 3551 controls. The outcome of interest is an indicator of survival at 30 days and there are 71 covariates selected by a panel of experts as variables that relate to practioners’ decisions about whether to use RHC (Connors et al. (1996)). We estimate the propensity score using a logistic regression with all covariates and age squared, as both younger and older patients are less likely to receive RHC. There is low to moderate overlap in the estimated propensity scores of the treated and control groups and many propensity scores are close to zero (Supplementary Figure 8); this indicates that there is potential for the standard ATE estimator to have large bias and variance. When applying IPW weights, the average absolute standardized mean difference (SMD) of treated and control covariates after weighting is $0.018$ and the maximum is $0.062$. Supplementary Figure 9 shows the 20 most imbalanced covariates after weighting, however in general the treated and control groups are relatively balanced after weighting. For illustrative purposes, we quantitatively assess treatment effect heterogeneity through a series of linear probability models and find minimal evidence that there is substantial treatment effect heterogeneity within the data (in practice this assessment would impact downstream inference). We apply our methods as described in Section 3 to characterize and select a set of optimal estimands. We use a grid of all combinations $c,d=0,0.05,0.10,\ldots,1$ and $1,000$ null distribution or bootstrap replications for all metrics.
Figure 6 shows the estimand mismatch and statistical bias $p$-value contours for this data. The points on the plot represent the set of optimal estimands selected by our proposed procedure. For the $p$-values pertaining to estimand mismatch, the estimands with $p>0.05$ generally follow a linear trend above the diagonal on the plot; for these estimands, $h(\bm{x})$ peaks around propensity scores of $0.38$, the proportion of treated individuals in the data. The $p$-values pertaining to our metric for characterizing statistical bias are largest for the estimands in the top right corner of the plot around the ATO. Of note, the statistical bias $p$-values are quite small (between 00 and $0.40$); only only estimands with $c,d\geq 0.8$ have $p$-values greater than $0.30$. This indicates many of the estimands have a high potential for statistical bias - especially for the large region of estimands that fall within the $[0,0.05]$ $p$-value contour. In comparison to the simulations, here we do not know the true propensity score model, and therefore the $p$-values here characterize the potential for statistical bias both from the lack of overlap and any propensity model misspecification.
Figure 6: Contours of the $p$-values corresponding to estimand mismatch and statistical bias for the right heart catheterization study. Points on the figures show the optimal estimands for each estimand mismatch $p$-value contour selected using our method in Section 3.3. Note that the estimand mismatch panel has contours of $[0,0.025]$ and $(0.025,0.05]$ while the statistical bias panel only has the contour $[0,0.05]$.
For the ATE, ATO, and the selected estimands in between them, we explore the tradeoffs between estimand mismatch, statistical bias, and variance as characterized by our three metrics (Figure 7). In Figure 7, each point represents a different estimand; points are positioned such that looking left to right, the estimands become less similar to the ATE and more similar to the ATO. Even though the energy distances are not directly comparable between estimands, the trends in these values can still be informative. The distance between the weighted covariate distributions and the sample covariate distribution (corresponding to estimand mismatch) is similar for most of the selected estimands, however starting around the $p=0.15$ estimand the distance starts to increase substantially to the right. The distance is largest at the ATO, where the estimand has maximal potential mismatch with the ATE. We see the opposite trend in the distance between the weighted treated and control propensity score distributions (corresponding to statistical bias); the ATE and similar estimands have the largest distance between these distributions. The distance then decreases as the estimands become more similar to the ATO, and the successive decrease between each estimand becomes less sharp around the $p=0.2$ estimand. The standard error of the estimators generally tends to decrease as the estimands become less similar to the ATE and more similar to the ATO.
Figure 7: Characteristics of the ATE, optimal selected estimands, and ATO for the right heart catheterization study. The selected estimands are labeled by the lower bound of their corresponding estimand mismatch $p$-value contour. Generally, lighter colors indicate estimands with better performance with respect to one aspect of the bias decomposition; lighter points indicate smaller potential for estimand mismatch, while lighter background shading indicates smaller potential for statistical bias.
There are multiple ways to select the single optimal estimand based on the tradeoffs between the three metrics. First, we reiterate that these estimands only differ when there is treatment effect heterogeneity. Therefore, in cases with minimal treatment efffect heterogeneity, estimand mismatch may be a lower priority than statistical bias regardless of whether the ATE population is of interest. For this data, there is moderate to low evidence of treatment effect heterogeneity, which indicates we may want to prioritize minimizing estimator bias and standard error rather than estimand mismatch. When focusing only on these two metrics, the $p=0$ estimand is the optimal estimand as it is in the largest statistical bias $p$-value contour and its estimator has the lowest estimated standard error. If we want to maximize the performance of our selected estimand while still maintaining a similar target population, we could choose either the $p=0.05$ or $p=0.10$ estimand as they, respectively, are in the highest p-value contour or have the lowest estimated standard error of all estimands with estimand mismatch $p>0.05$. If there is belief in or evidence of substantial treatment effect heterogeneity, we would potentially want to select an estimand with smaller estimand mismatch that also has some statistical bias and variance benefits, such as the $p=0.20$ estimand. For this data, any of the $p=0$, $p=0.05$, or $p=0.10$ estimands may be the optimal choice due to the limited treatment effect heterogeneity exhibited in the data.
While in practice one would not look at the actual treatment effect estimates before selecting an estimand, for illustrative purposes we show the different treatment effect estimates for the set of optimal estimands in the bottom right panel of Figure 7. The treatment effect estimate is around $-0.056$ for the ATE estimator and tends to decrease as the estimands become more similar to the ATO, whose estimate is around $-0.065$; all estimates indicate a negative effect of RHC on 30 day survival. The ATO and all estimands from estimand mismatch contours with $p<0.3$ have similar treatment effect estimates around $-0.065$. The ATE estimator has potential for substantial bias, as the statistical bias $p$-value is less than $0.05$ and further has large variance. A re-analysis of this data by Hirano and Imbens (2001) using regression augmented estimators (which are more robust to model misspecification) found that when using propensity score and regression models with at least $15$ covariates, the treatment effect estimates were around $-0.062$ (ranging from $-0.059$ to $-0.067$). Since these results are similar to treatment effect estimates from the ATO and similar estimands, there is evidence that these estimands may yield estimates closer to the true ATE than the ATE estimator due to a reduction statistical bias while not incurring substantial estimand mismatch.
Discussion
Covariate imbalance and a lack of overlap pose significant challenges when estimating a causal treatment effect in a target population. To address these challenges, we have introduced a framework for characterizing and selecting causal estimands based on the population targeted and the corresponding estimator’s statistical bias and variance. We have proposed a bias decomposition and two design-based metrics that characterize bias due to estimand mismatch (i.e., deviation from the population of interest) and statistical bias, allowing one to characterize how the choice of estimand impacts performance along these facets. Our estimand selection procedure identifies a sequence of optimal estimands with a clear interpretation in respect to the bias-variance tradeoff of their corresponding estimators. In addition, our selection procedure is flexible and allows the analyst to incorporate their domain-specific preferences for maintaining the target population versus reducing statistical bias and variance. When applying our method to simulated data with a population target estimand of the ATE, the estimands selected through our procedure reduce estimator MSE by moderating the size of both bias and variance, rather than minimizing one of these attributes.
While we present methods and results with respect to the ATE as the population target estimand, these methods can be applied to any other estimand such as the ATT or when transporting effects to a new population. In addition, our methods could be extended to regression augmented estimators, though further work is needed to assess how well our proposed metric (Section 3.2) characterizes statistical bias in these cases. Furthermore, while we explore the set of estimands in Equation (4), any potential set of estimands defined by weighted populations could be explored; for example, estimands with $c,d<0$ or $c,d>1$ could be of interest, especially when the target estimand is the ATT or ATC. Of note, identifying a different set of potential estimands may be important when transporting effects to a new population as the there is no guarantee that a new population is similar to the weighted populations defined by Equation (4).
We acknowledge that using our proposed procedure to identify an estimand could influence downstream inference. However, our procedure is design-based which helps protect against potential impacts on inference when using an estimand selected through this procedure. In addition, Yang and Ding (2018) provide methods for incorporating uncertainty from both the design and analysis stage in the inferential procedure when using continuous weights. Since all potential weights from Equation (4) are continuous, these methods could be used here.
While this work is presented as a unified framework for estimand selection in the presence of a lack of overlap, the proposed metrics could be useful in a variety of different scenarios. For example, since our metric for characterizing statistical bias is sensitive to bias due to propensity score model misspecification, it could be used to assess the fit of propensity score models in relation to a single estimand. In addition, variations of our metric for characterizing estimand mismatch could be useful when comparing a matched sample to the original sample population or when identifying an optimal target population when generalizing or transporting effects.
Funding
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2237827.
Supplementary Materials
Proof of Proposition 1: By Theorem 2 of Li, Morgan, and Zaslavsky (2018), for a given sample $\bm{X}=\{x_{1},\ldots,x_{p}\}$
[…]
For clarity in notation, $E[\cdot]$ indicates $\int\cdot f(\bm{x})d\bm{x}$. Using the same argument as the proof of Corollary 1 in Li, Morgan, and Zaslavsky (2018), by the Cauchy-Schwartz inequality
$E\{h(\bm{x})\}^{2}$ $=\left[E\left\{\frac{h(\bm{x})}{\sqrt{h^{*}(\bm{x})}}\sqrt{h^{*}(\bm{x})}\right\}\right]^{2}$
$\leq E\left\{\frac{h(\bm{x})^{2}}{h^{*}(\bm{x})}\right\}E\{h^{*}(\bm{x})\}$
where equality holds when $h(\bm{x})\propto h^{*}(\bm{x})$. Applying this result to the right-hand side of (12), we have that
[…]
and $h(\bm{x})\propto h^{*}(\bm{x})$ gives the smallest asymptotic variance for $\hat{\tau}_{h}$ among all $h$.
Figure 1: Propensity score overlap for all simulation scenarios. Panel a) and b) show overlap for simulated data with 25% and 50% of individuals treated, respectively. Note that $\gamma$ indicates the level of propensity score overlap where $\gamma=1$ is high overlap and $\gamma=4$ is low overlap.
Figure 2: The left panel presents the average of the minimum of the two estimand mismatch related $p$-values, calculated as described in Section 3.1, for each estimand. The right panel presents the absolute difference between the true estimand values and the true ATE (i.e., the absolute estimand mismatch) under the medium $\Delta$ heterogeneity scenario when simulating the treatment effect as a linear function of the propensity score.
Figure 3: Proportion of estimands within each statistical bias associated $p$-value contour with absolute percent bias greater than 1%, 5%, 10%, and 25%.
Figure 4: Average bootstrapped standard error for all estimators corresponding to the estimands defined by Equation (4) across all simulation scenarios. Each of the 16 subplots corresponds to a single simulation scenario as labelled by the top and right axes.
Figure 5: Spearman correlation between the true estimator standard error (as measured by the empirical standard deviation) and the average bootstrapped standard error across all simulation scenarios. Each of the 16 subplots corresponds to a single simulation scenario as labelled by the top and right axes.
Figure 6: Mean squared error (MSE) with respect to the true ATE for the estimators for the ATE, the estimands selected with our procedure, and the ATO by simulation scenario. Each of the 16 subplots corresponds to a single simulation scenario as labelled by the top and right axes. Estimands are labeled by the lower bound of their corresponding estimand mismatch $p$-value contour. Estimands selected (i.e., estimand mismatch $p$-value contours present) in at least 900 of the simulated datasets are shown.
Figure 7: Average of the selected estimands (i.e., average of $c$ and $d$) for each estimand mismatch $p$-value contour across all possible simulated datasets for all simulation scenarios. Each of the 16 subplots corresponds to a single simulation scenario as labelled by the top and right axes. Estimands are labeled by the lower bound of their corresponding estimand mismatch $p$-value contour.
Figure 8: Estimated propensity score distributions of treated and control groups for the right heart catheterization study (Connors et al., 1996).
Figure 9: Twenty most imbalanced covariates after IPW weighting for the right heart catheterization study (Connors et al., 1996).